OpenStudy (shadyykatie):

1. List the sides ∆ABC in order from least to greatest. A. AC, BC, AB B. AC, AB, BC C. AB, AC, BC D. AB, BC, AC 2. Which of the following are NOT the lengths of a triangle? A. 2, 3, 4 B. 2, 3, 2 C. 2, 3, 3 D. 2, 3, 6 3. The Polygon Angle-Sum states: The sum of the measures of the angles of an n-gon is ______? A. n - 2/180 B. (n - 1)/180 C. 180/n - 1 D. (n - 2)/180

OpenStudy (shadyykatie):

@arabpride

OpenStudy (shadyykatie):

# 1

OpenStudy (shadyykatie):

@arabpride

OpenStudy (campbell_st):

find the size of angle C use angle sum of a triangle.. then use the rule the longest side in a triangle is opposite the largest angle.. the shortest side is opposite the smallest angle... middle length opposite the middle sized angle.. hope it helps

OpenStudy (campbell_st):

for question 2 the sum of the 2 shorter sides must be greater than the measure of the longest side.

OpenStudy (shadyykatie):

I am confussed :( sorry I am not good at this

OpenStudy (anonymous):

Q1 ans CB is the largest side then comes AC and AB is the smallest the side opposite to the largest angle is the largest side. you can find angle C bu using the angle sum property! angleA + angleB + angleC = 180 degrees 75 + 58 + angleC = 180 angle C= 180-133 = 47 degrees!! good luck!!

OpenStudy (anonymous):

that is option C AB,AC,BC